OpenStudy (anonymous):

Solve: 3(x+2)=2(2x-1)-12

OpenStudy (anonymous):

3x+6=4x-2-12.........-x=-2-12-6.......x=20

OpenStudy (turingtest):

yup

OpenStudy (anonymous):

You can use distributive property and get: x=20 ^_^